# GreekBarBench: A Challenging Benchmark for Free-Text Legal Reasoning and Citations

### v1

@@ 1 Introduction
All resources including the two benchmarks (except for a small semi-private test set) and the prompts are publicly available. [^3] The URL of the GitHub repository will be provided in the camera-ready version.
@@ 2.2 Task
Benchmark Lang Citations Multi-Hop Free-Text Judge Eval
@@
Table 2: Comparison of legal benchmarks. GreekBarBench uniquely encompasses all challenging features essential for evaluation in realistic and practical scenarios. ‘Lang’: language of dataset. ‘Citations’: legal articles. ‘Multi-hop’: reasoning using multiple sources. ‘Free-text’: open-ended responses. ‘Judge Eval’: manual evaluations to compare LLM-judges.
@@ 2.3 Dataset Statistics
Answering exam questions requires citing articles from 25 legal code documents, which we collect from the same source website as the exams.666www.lawspot.gr Detailed statistics for these documents are presented in Table 3 per legal area. Articles are cited 931 times in total, across all exam questions. The articles within each legal code document are grouped thematically into chapters. The total number of citable articles is more than 16 thousand.
@@ 5.1 Models
Our experiments evaluate a diverse range of LLMs, comprising proprietary models (from OpenAI, Google, Anthropic) and open-weight models; Deepseek-R1 (DeepSeek-AI et al., 2025), Gemma-3 (Team et al., 2025), and Llama-Krikri-8B [^8] https://huggingface.co/ilsp/Llama-Krikri-8B-Instruct, a model specifically pretrained for the Greek language. We accessed proprietary models and the large open-weight Deepseek-R1 through Application Programming Interfaces (APIs) provided by OpenAI, Google, and AWS. The remaining open-weight models were deployed on a cluster of eight A100 GPUs using the vLLM framework (Kwon et al., 2023). Due to limited resources, we only evaluate a single run for each model. We used the default parameter configurations as specified by each model’s provider.
@@ 5.2 Manual Evaluation by Legal Experts
[…]
L-Krikri-8B 0.747 0.751 -
Gemma-3-27B 0.819 0.749 -
[…]
@@ 6 Related Work
In the legal domain, LexGLUE (Chalkidis et al., 2022) and LEXTREME (Niklaus et al., 2023) are established benchmarks for legal classification tasks. LegalBench (Guha et al., 2023) is the standard for evaluating LLMs on legal reasoning via multiple-choice questions. More closely related to our work, task 5.4 of LexEval (Li et al., 2024) uses a similar legal examination for Chinese instead of Greek, but, unlike our approach, LexEval does not provide citations or use LLM-as-a-judge, instead evaluating with the less reliable, overlap-based ROUGE metric (Cohan and Goharian, 2016). LLeQA (Louis et al., 2024) collects pairs of everyday legal questions and answers, including citations to French statutory law articles, but they evaluate using the METEOR metric without measuring its correlation with human experts. CaseGen (Li et al., 2025) on the other hand, assesses document drafting and legal judgment generation in Chinese using the LLM-as-a-judge approach. While they do measure agreement between human and LLM evaluations, they do not compare different prompts or models. Concurrent work, OAB-Bench (Pires et al., 2025), uses data from the Brazilian Bar Examination and also provides the official guidelines as rubrics for the LLM-judges. They do evaluate different LLM-judges, but they only do it for three samples and they do not provide citations to statutory articles. Notably, the complexity of their rubrics necessitates the use of the expensive OpenAI-o1 model for evaluation, thus significantly increasing the overall cost, amounting to approximately \$50 for each LLM evaluated.
@@
Important Natural Language Processing resources for the Greek language include classification models (Koutsikakis et al., 2020; Saketos et al., 2024), alongside more recent LLMs pretrained on Greek like Meltemi [^11] Meltemi was excluded from our experiments, because of its relatively small context length of 8 billion tokens. (Voukoutis et al., 2024) and Llama-Krikri [^12] https://huggingface.co/ilsp/Llama-Krikri-8B-Instruct, which we tested in our experiments (§ 5.4). Existing Greek legal datasets cover only classification and summarization tasks (Angelidis et al., 2018; Papaloukas et al., 2021; Koniaris et al., 2023). Although Greek LLM benchmarks exist for other domains, such as finance (Peng et al., 2025) and medicine (Voukoutis et al., 2024), the legal domain currently lacks one.

### v2

@@ 1 Introduction
All resources including code, the two benchmarks (except for a small semi-private test set) and the prompts are publicly available. [^3] Code and prompts: https://github.com/nlpaueb/greek-bar-bench
Dataset: https://huggingface.co/datasets/AUEB-NLP/greek-bar-bench
@@ 2.2 Task
Benchmark Lang Citations Multi-Hop Free-Text Judge-Eval
@@
Table 2: Comparison of legal benchmarks. GreekBarBench uniquely encompasses all challenging features essential for evaluation in realistic and practical scenarios. ‘Lang’: language of dataset. ‘Citations’: legal articles. ‘Multi-hop’: reasoning using multiple sources. ‘Free-text’: open-ended responses. ‘Judge-Eval’: separate benchmark to compare LLM-judges (‘-’ means that the benchmark exists but is not publicly available).
@@ 2.3 Dataset Statistics
Answering the exam questions requires citing articles from 25 legal code documents, obtained from the official website of the Greek government. [^6] https://search.et.gr/el/ Detailed statistics for these documents are presented in Table 3 per legal area. Articles are cited 931 times in total, across all exam questions. The articles within each legal code document are grouped thematically into chapters. The total number of citable articles is more than 16 thousand.
@@ 5.1 Models
Our experiments evaluate a diverse range of LLMs, comprising proprietary models (from OpenAI, Google, Anthropic) and open-weight models; Deepseek-R1 (DeepSeek-AI et al., 2025), Gemma-3 (Team et al., 2025), and Llama-Krikri-8B (Roussis et al., 2025); a model specifically pretrained for the Greek language. We accessed proprietary models and the large open-weight Deepseek-R1 through Application Programming Interfaces (APIs) provided by OpenAI, Google, and AWS. The remaining open-weight models were deployed on a cluster of eight A100 GPUs using the vLLM framework (Kwon et al., 2023). Due to limited resources, we only evaluate a single run for each model. We used the default parameter configurations as specified by each model’s provider.
@@ 5.2 Manual Evaluation by Legal Experts
GPT-4.1-mini 0.723 0.856 \$\$
GPT-4.1 0.807 0.855 \$\$\$
Gemini-2.0-F 0.751 0.794 \$
L-Krikri-8B 0.747 0.751
Gemma-3-27B 0.819 0.749
Gemini-2.0-L-F 0.695 0.708 \$
GPT-4.1-nano 0.542 0.372 \$
Table 4: Comparison of judges on GBB-JME, using Simple-Judge (Simple-J) and Span-Judge (Span-J). The cost per one million input tokens is denoted by: \$ (less than \$0.3), \$\$ (less than \$1), and \$\$\$ (less than \$3).
@@ 6 Related Work
In the legal domain, LexGLUE (Chalkidis et al., 2022) and LEXTREME (Niklaus et al., 2023) are established benchmarks for legal classification tasks. LegalBench (Guha et al., 2023) is the standard for evaluating LLMs on legal reasoning via multiple-choice questions. More closely related to our work, task 5.4 of LexEval (Li et al., 2024) is based on Chinese Bar exams, but, unlike our approach, it does not provide citations or use LLM-as-a-judge, instead evaluating with the less reliable, overlap-based ROUGE metric (Cohan and Goharian, 2016). Similarly, LLeQA (Louis et al., 2024) contains everyday legal questions in French, but evaluation is based on the METEOR metric without measuring its correlation with human experts. CaseGen (Li et al., 2025) assesses document drafting and legal judgment generation in Chinese using the LLM-as-a-judge approach.
Two concurrent works, OAB-Bench (Pires et al., 2025) and LEXam (Fan et al., 2025), are most similar to GreekBarBench as they also provide open-ended legal questions and rubrics for LLM-judges. However, both have notable limitations. OAB-Bench, based on the Brazilian Bar exams, uses official guidelines as rubrics, but their complexity necessitates a costly model (OpenAI-o1), resulting in an evaluation cost of approximately \$50 per LLM. LEXam is based on English and German law school questions, which, unlike Bar exams, are targeted at less experienced students rather than graduates. Crucially, while both projects evaluate the quality of LLM-judges, neither provides these manual evaluations as a publicly available, standalone benchmark, preventing the community from using them to assess new judges in the future.
@@
Important Natural Language Processing resources for the Greek language include classification models (Koutsikakis et al., 2020; Saketos et al., 2024), alongside more recent LLMs pretrained on Greek like Meltemi [^10] Meltemi was excluded from our experiments, because of its relatively small context length of 8 billion tokens. (Voukoutis et al., 2024) and Llama-Krikri-8B (Roussis et al., 2025), which we tested in our experiments (§ 5.4). Existing Greek legal datasets cover only classification and summarization tasks (Angelidis et al., 2018; Papaloukas et al., 2021; Koniaris et al., 2023). Although Greek LLM benchmarks exist for other domains, such as finance (Peng et al., 2025) and medicine (Voukoutis et al., 2024), the legal domain currently lacks one.